1. When a man stands, his brain is 0.5 m above his heart. When he bends, his brain is 0.4m below his heart, by how much does his blood pressure in his brain change?
  2. At what acceleration would you expect the blood pressure in the brain to drop to zero in an erect person ? ( assume there are no body mechanism operating to compensate for such conditions )
1
Expert's answer
2021-07-05T08:43:42-0400

1)


"\Delta p=\rho g h\\\Delta p=(1060)(9.8)(0.5-0.4)=1040\ Pa"

2)


"1040=(1060)(2)a\\a=0.49\frac{m}{s^2}"